# Grounding Natural Language to SQL Translation with Data-Based Self-Explanations

### v1

@@ IV-D Explanation as NL2SQL Verifier
Once the NL explanation for a specific query result is generated, the subsequent question for Cyclesql is how to effectively leverage the NL explanation to validate the translation. Drawing inspiration from the recent achievements in the NLP field, we thus formulate the translation validation problem as a textual entailment task (a.k.a. NLI task) [56, 57] and employ deep learning techniques to construct an NLI model for this purpose. Here, the NLI model aims to evaluate the semantic coherence between a given NL query and the NL explanation associated with a query result. The NL query fragment, labeled as hypothesis ($H$), is inferred from the NL explanation fragment known as premise ($P$) [^7] To provide extra context, we also include the query result along with the SQL query within the premise, using a special token (||||) to separate each part.. In other words, an explanation $P$ entails an NL query $H$, if $H$ is considered to be true within the context of the corresponding text $P$.
@@
Figure 7 presents the architecture of the dedicated NLI model. Building upon prior works [61, 62], we adopt a similar approach to construct the model, refining it via fine-tuning for specific classification objectives. We use a unified text encoder to jointly encode the hypothesis and the premise, applying a pooling layer to create a unified embedding vector, which is then fed into a classification layer for output.
@@ V-A1 Experimental Setup
  • Spider [7] includes $10,181$ NL queries and $5,693$ unique SQL queries on $206$ databases covering $138$ domains. As it uses different databases across its train, dev, and test data sets, the model’s generalizability can be properly evaluated.
@@
Baseline NL2SQL Models. We consider the following two types of baselines: Seq2seq-based translation models fine-tuned on the Spider training set, including SmBoP [14], Picard [16], and Resdsql [17], and LLM-based models, such as Gpt-3.5-turbo, Gpt-4, Chess [32], and Dailsql [22]. We briefly describe each baseline model as follows:
@@
  • Resdsql represents the question and schema in a tagged sequence. Resdsql was implemented using three scales of T5 model: Base, Large and 3B. In our experiments, we use Resdsql with the latter two scales, denoted to as $\textsc{Resdsql}_{\textsc{Large}}$ and $\textsc{Resdsql}_{\textsc{3b}}$, and exclude its combination with Natsql [66].
@@ V-A2 Results on Spider
Table 8(a) shows that upon integration with Cyclesql, most baselines can find the “best” translation results within $1\sim 2$ iterations, except Picard, which may require about $4$ iterations due to relatively lower quality of its sampled SQL queries as we mentioned earlier. On the other hand, as depicted in Figure 8(b), given that each baseline model can generate a list of top-k query candidates in a single inference process, the additional execution time overhead incurred by Cyclesql is minimal [^13] Due to the decoding process of Picard necessitating interaction with a web service to validate each predicted token, the inference time of Picard is time-consuming. Therefore, we omit it from this comparison.. This showcases that Cyclesql strikes a favorable balance between translation and inference latency in its current settings. In addition, the inference time of different models heavily depends on their respective model sizes. For example, SmboP with approximately $360$ million parameters, demonstrates notably faster performance compared to $\textsc{Resdsql}_{\textsc{3b}}$, which boasts around $3$ billion parameters.
@@ V-A4 Analysis of Cyclesql
Figure 9 presents the EX results across the four Spider-related benchmarks. It is evident that relying on the feedback generated by the SQL2NL approach becomes unreliable for the models, which may even lead to a negative impact on their performance when applying to $\textsc{Resdsql}_{\textsc{Large}}$. The findings emphasize the importance of feedback quality, demonstrating that the feedback generated by the designed approach used in Cyclesql embeds rich semantics, thus establishing a reliable feedback loop inside the NL2SQL process.
@@ V-B Qualitative Evaluation
Results and Analysis. The feedback is positive for Cyclesql: $14$ out of $20$ participants agreed that the explanations generated by Cyclesql are more comprehensible to them compared with the ones from Gpt-3.5-turbo. Moreover, the average rating scores consistently surpass those of Gpt-3.5-turbo, as illustrated in Figure 10(b). More importantly, participants confirmed that examining these explanations can greatly assist them in understanding the reasons behind the query results and may enable them to detect potential translation errors that happened with the underlying NL2SQL process.
@@ VI Related Work
Natural Language Interface for Database. NLIDBs have been studied both in the database management and NLP communities. Many machine learning-based approaches treat the NL2SQL problem as a Seq2seq translation task and leverage the encoder-decoder architecture to do the translation [6, 7, 12, 11, 13, 73, 74, 14, 15, 17, 67, 75, 76]. With the excellent performance of LLMs in various NLP tasks, recent works have explored applying LLMs to the NL2SQL task [40, 41, 19, 77, 20], mostly in an end-to-end manner.
@@ VII Conclusion & Future Work
In this paper, we proposed an iterative framework named Cyclesql for existing end-to-end models to autonomously generate the best output using NL explanations as self-provided feedback. Cyclesql initials by tracking data provenance information and query semantics for a given query result and utilizes text generation technique to generate an NL explanation. The NL explanation is then used to validate the correctness of the underlying NL2SQL translation, ultimately resulting in a better translation outcome. We conducted extensive experiments to study the effectiveness of Cyclesql on five public benchmarks. The results showed Cyclesql could consistently improve the translation performance of the existing seven end-to-end models on these benchmarks. Moreover, we conduct additional qualitative experiments with a case study and a user study to analyze the quality of the explanations, which gain more insights on Cyclesql.
While Cyclesql has demonstrated its effectiveness in its current form, these findings underscore the necessity for further future efforts in this direction. One potential area of investigation is how to close the feedback loop with human involvement, enabling Cyclesql to better align with humans and adapt its translation over time. In addition, developing a more reliable way to assist data tracking in Cyclesql is critical for Cyclesql, especially in better supporting empty-result queries. Finally, an intended future research direction involves exploring an effective way to incorporate fine-grained semantics both in NL queries and explanations during the training of the NLI model, thereby enhancing its resilience in managing comparable translations.

### v2

@@ IV-D Explanation as NL2SQL Verifier
Once the NL explanation for a specific query result is generated, the subsequent question for Cyclesql is how to effectively leverage the NL explanation to validate the translation. Drawing inspiration from recent achievements in the NLP field, we thus formulate the translation validation problem as a textual entailment task (a.k.a. NLI task) [56, 57] and employ deep learning techniques to construct an NLI model for this purpose. Here, the NLI model aims to evaluate the semantic coherence between a given NL query and the NL explanation associated with a query result. The NL query fragment, labeled as hypothesis ($H$), is inferred from the NL explanation fragment known as premise ($P$) [^7] To provide extra context, we also include the query result along with the SQL query within the premise, using a special token (||||) to separate each part.. In other words, an explanation $P$ entails an NL query $H$, if $H$ is considered to be true within the context of the corresponding text $P$.
@@
Fig. 7 presents the architecture of the dedicated NLI model. Building upon prior works [61, 62], we adopt a similar approach to construct the model, refining it via fine-tuning for specific classification objectives. We use a unified text encoder to jointly encode the hypothesis and the premise, applying a pooling layer to create a unified embedding vector, which is then fed into a classification layer for output.
@@ V-A1 Experimental Setup
  • Spider [7] includes $10,181$ NL queries and $5,693$ unique SQL queries of $206$ databases covering $138$ domains. As it uses different databases across its train, dev, and test data sets, the model’s generalizability can be properly evaluated.
@@
Baseline NL2SQL Models. We consider the following two types of baselines: Seq2seq-based translation models fine-tuned on the Spider training set, including SmBoP [14], Picard [16], and Resdsql [17], and LLM-based models, such as Gpt-3.5-turbo, Gpt-4, Chess [32], and Dailsql [22]. We briefly describe each baseline model as follows.
@@
  • Resdsql encodes the question and schema as a tagged sequence and is implemented with T5 model in Base, Large and 3B scales. Our experiments use Resdsql with the latter two scales, denoted to as $\textsc{Resdsql}_{\textsc{Large}}$ and $\textsc{Resdsql}_{\textsc{3b}}$, and exclude its combination with Natsql [66].
@@ V-A2 Results on Spider
Table 8(a) shows that upon integration with Cyclesql, most baselines can find the “best” translation results within $1\sim 2$ iterations, except Picard, which may require about $4$ iterations due to relatively lower quality of its sampled SQL queries as we mentioned earlier. On the other hand, as depicted in Fig. 8(b), given that each baseline model can generate a list of top-K query candidates in a single inference process, the additional execution time overhead incurred by Cyclesql is minimal [^13] Due to the decoding process of Picard necessitating interaction with a web service to validate each predicted token, the inference time of Picard is time-consuming. Therefore, we omit it from this comparison.. This showcases that Cyclesql strikes a favorable balance between translation and inference latency in its current settings. In addition, the inference time of different models heavily depends on their respective model sizes. For example, SmboP with approximately $360$ million parameters, demonstrates notably faster performance compared to $\textsc{Resdsql}_{\textsc{3b}}$, which boasts around $3$ billion parameters.
@@ V-A4 Analysis of Cyclesql
Fig. 9 presents the EX results across the four Spider-related benchmarks. It is evident that relying on the feedback generated by the SQL2NL approach becomes unreliable for the models, which may even lead to a negative impact on their performance when applying to $\textsc{Resdsql}_{\textsc{Large}}$. The findings emphasize the importance of feedback quality, demonstrating that the feedback generated by the designed approach used in Cyclesql embeds rich semantics, thus establishing a reliable feedback loop inside the NL2SQL process.
@@ V-B Qualitative Evaluation
Results and Analysis. The feedback is positive for Cyclesql: $14$ out of $20$ participants agreed that the explanations generated by Cyclesql are more comprehensible to them compared with the ones from Gpt-3.5-turbo. Moreover, the average rating scores consistently surpass those of Gpt-3.5-turbo, as illustrated in Fig. 10(b). More importantly, participants confirmed that examining these explanations can greatly assist them in understanding the reasons behind the query results and may enable them to detect potential translation errors that happened with the underlying NL2SQL process.
@@ VI Related Work
Natural Language Interface for Database. NLIDBs have been studied both in database management and NLP communities. Many machine learning-based approaches treat the NL2SQL problem as a Seq2seq translation task and leverage the encoder-decoder architecture to do the translation [6, 7, 12, 11, 13, 73, 74, 14, 15, 17, 67, 75, 76, 77]. With the excellent performance of LLMs in various NLP tasks, recent works have explored applying LLMs to the NL2SQL task [40, 41, 19, 78, 20, 79], mostly in an end-to-end manner.
@@ VII Conclusion & Future Work
In this paper, we proposed an iterative framework named Cyclesql for end-to-end models to autonomously generate the best output using NL explanations. Cyclesql initials by tracking data provenance and query semantics for a given query result to generate an NL explanation. The explanation is then used to validate the correctness of the underlying translation, ultimately resulting in a better translation outcome. We conducted extensive experiments on five benchmarks, showing that Cyclesql consistently improves translation performance across seven models. Additional case and user studies provided further insights into the quality of the generated explanations.
While Cyclesql has proven effective, further efforts are needed to enhance its capabilities. One potential area of investigation is how to close the feedback loop with human involvement, enabling Cyclesql to better align with humans and adapt its translation over time. In addition, developing a more reliable way to assist data tracking in Cyclesql is critical for Cyclesql, especially in better supporting empty-result queries. Finally, an intended future research direction involves exploring an effective way to incorporate fine-grained semantics both in NL queries and explanations during the training of the NLI model, thereby enhancing its resilience in managing comparable translations.
VIII Acknowledgments
The authors would like to thank all the anonymous reviewers for their insightful comments and suggestions. This work was supported by NSFC (Grant No. 62272106).